OpenStudy (anonymous):

The graph of the following system of equations is 2x + y = 6 6x + 3y = 18 A) overlapping lines B) parallel lines C)intersecting lines can someone help me please

OpenStudy (anonymous):

I think its c but im not sure

OpenStudy (abmon98):

Parallel lines share the same gradient so convert them into slope intercept form y=-2x+6 3y=-6x+18 Divide by 3 y=-2x+6 Parallel lines never intersect

OpenStudy (anonymous):

oh so it is parallel?

OpenStudy (aum):

Divide the second equation throughout by 3. What do you get?

OpenStudy (aum):

6x + 3y = 18 divide by 3 2x + y = 6 The second equation is identical to the first equation. Therefore, the two equations represent one and the same line. They are overlapping lines.

OpenStudy (anonymous):

so it is a

OpenStudy (abmon98):

yes @aum is right am sorry they are both the same line

OpenStudy (anonymous):

ok thanks for the help